OpenStudy (anonymous):

find the equation on the line containing the given pair of points express your answers in the form x=a,y=b or y=mx+b what is the equation of the line

OpenStudy (saifoo.khan):

b = am + b

OpenStudy (anonymous):

(-4,-4) and (1,1) so y=?

OpenStudy (saifoo.khan):

y intercept?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what is the equation of those so y=

OpenStudy (saifoo.khan):

fine the slope first.

OpenStudy (saifoo.khan):

slope is 1

OpenStudy (anonymous):

ok

OpenStudy (saifoo.khan):

fint the y intercept now!

OpenStudy (saifoo.khan):

find*

OpenStudy (saifoo.khan):

(1) = 1(1) + b b = 0

OpenStudy (saifoo.khan):

therefore: y = mx

OpenStudy (anonymous):

huh im so lost

OpenStudy (saifoo.khan):

y?

OpenStudy (anonymous):

how did you get that

OpenStudy (saifoo.khan):

u got a picture of a heroine.

OpenStudy (saifoo.khan):

Let me give u a shortcut.

OpenStudy (saifoo.khan):

(x2 - x1) = (y2 - y1)/m